A sample of 16 ten-year-old girls had a mean weight of 71.5 and a standard deviation of 

12 pounds, respectively. Assuming normality, find the 90, 95, and 99 percent confidence 

intervals forthe mean of the population.


1
Expert's answer
2021-08-10T11:31:25-0400

"CI=\bar {X}\pm t_{\frac{\alpha}{2},n-1}\frac{s}{\sqrt{n}}"

90%CI

"CI=\bar {X}\pm t_{0.05,n-1}\frac{s}{\sqrt{n}}"

"\bar{X}=71.5"

"s=12"

"n=16"

"t_{0.05,15}=1.753"

"CI=71.5\pm 1.753(\frac{12}{\sqrt{16}})"

=[66.242, 76.759]


95%CI

"CI=\bar {X}\pm t_{0.025,n-1}\frac{s}{\sqrt{n}}"

"t_{0.025,15}=2.131"

"CI=71.5\pm 2.131(\frac{12}{\sqrt{16}})"

"=[65.106, 77.894]"


99%CI

"CI=\bar {X}\pm t_{0.005,n-1}\frac{s}{\sqrt{n}}"

"t_{0.005,15}=2.947"

"CI=71.5\pm 2.947(\frac{12}{\sqrt{16}})"

"=[62.66, 80.34]"